Circular No. 9119

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

V2674 OPHIUCHI = NOVA OPHIUCHI 2010 No. 2 As first announced on CBET 2176, S. Nakano (Sumoto, Japan) re-

ported the discovery by H. Nishimura (Miyawaki, Kakegawa, Shizuoka-ken, Japan) of a possible nova on frames taken on Feb. 18.845 UT using a Minolta 120-mm f/3.5 lens and a Canon EOS 5D Digital camera mounted on a Takahashi EM100 equatorial telescope; Nakano measured mag 9.4 for the variable from Nishimura’s JPEG image (limiting mag 12.7), along with the following precise position (uncertainty ± 4′′): α = 17^{h}26^{m}32^{s}.19, δ = −28^{o}49′36′′.3 (equinox 2000.0). Nishimura notes that nothing is visible at this position on his frames taken during 2009 Feb.-2010 Feb. 13.84 (limiting mag 11). Additional CCD magnitudes (unfiltered unless noted otherwise) and position end figures for the nova: Feb. 19.48, R_{c} = 9.02, 32^{s}.11, 38′′.5 (S. Kiyota, Tsukuba, Japan, 0.30-m reflector, remotely from near Mayhill, NM, U.S.A.); 19.778, 9.0, 32^{s}.08, 38′′.5 (K. Itagaki, Takanezawa station, Tochigiken, Japan, 0.30-m reflector); 19.825, 8.9, 32^{s}.11, 39′′.0 (K. Kadota, Ageo, Japan, 0.25-m reflector); 20.5, 9.4:, 32^{s}.15, 41′′.8 (E. Guido and G. Sostero, 0.25-m reflector at the GRAS Observatory near Mayhill, NM, U.S.A.). Kiyota adds that nothing is visible at this position on several Digitized Sky Survey plates. K. Imamura and K. Tanabe, Okayama University of Science, report (via H. Yamaoka, Kyushu University) that a low-resolution spectrum of the presumed nova, obtained on Feb. 19.850 UT with a 0.28-m telescope, shows a clear Hα feature with a P-Cyg profile, which suggests that the object is a genuine classical nova; an absorption minimum of the Hα feature is ∼ 1000 km/s, blueshifted from the emission peak. M. Fujii, Kurashiki, Okayama, Japan, writes (also via Yamaoka) that a low-resolution spectrum of this object, obtained on Feb. 19.86 UT with his 0.28-m telescope, shows — in addition to the Balmer lines with P-Cyg profile — a Fe II (37, 42, 49) emission feature, which suggests that the variable is a “Fe II”-class nova. Fujii measured that the blueshift of the Hα absorption minimum from the emission peak is ∼ 700 km/s, and the equivalent width of Na I D lines is ∼ 0.4 nm. Additional details concerning this nova can be found on CBETs 2176, 2179, and 2185.

E. Kazarovets and N. N. Samus report that the formal GCVS designation V2674 Oph has been assigned to this nova.

COMET 233P/LA SAGRA Following an identification by G. V. Williams of comet P/2009 WJ_{50}

(cf. IAUC 9117) with 2005 JR_{71} (cf. MPO 169487; MPEC 2010-D2; MPC 68903), it has been given the permanent number 233P (cf. MPC 68628).

2010 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green